# Data Selection via Optimal Control for Language Models

### v1

@@ 2.3.2 Data Scorer
where $x_{n}^{\mathrm{prx}}$ is the $n^{\text{th}}$ instance in $\mathcal{D}^{\mathrm{prx}}$, $\bm{\phi}$ is the parameters of the data scorer, and $\overline{\bm{h}}(\cdot,\cdot)\in\mathbb{R}^{d}$ is the average output hidden states of an LM along the sequence length, with $d$ representing the hidden state size. ${\bm{w}}\in\mathbb{R}^{d},b\in\mathbb{R}$ are the parameters of the linear head. After fine-tuning, we infer the data quality scores of the instances in $\mathcal{D}$ with the data scorer, where the quality score for $x_{n}$ is given by $\gamma(x_{n})={{\bm{w}}^{*}}^{\top}\overline{{\bm{h}}}(x_{n},\bm{\phi^{*}})+b^{*}$.
@@ 2.4 Discussion
Compared to existing offline data selection methods that curate the pre-training corpus before the LM training starts using pattern information [79], deduplication [70, 1], or single-step checkpoints [22], PDS incorporates long-range training dynamics into data selection, as reflected by the “target vector” ${\bm{\lambda}}^{*}_{t}$ in Eq. (5). This can be critical for selecting high-quality instances, given the highly dynamic nature of LM pre-training. Although we run Algorithm 1 in a proxy environment and transfer the quality scores to the large-scale setting via the data scorer, many previous works [80, 81] have shown that data quality information is learnable and transferable across model sizes. Different LMs also share many common training dynamics [69].
@@
Unlike recent online data selection approaches to incorporate LM training dynamics [81, 75], PDS selects the pre-training corpus offline. This allows PDS to be run only once and used for pre-training multiple LMs of any sizes, without incurring additional computational overhead. The FLOPs needed by PDS in a proxy environment are also negligible compared to the demands of large-scale pre-training, as shown in a complexity analysis in Section 3.3. Besides, the offline nature of PDS makes it flexible to be integrated into optimized pre-training pipelines [15] by simply replacing the data sources.
@@ 3.1 Experimental Setup
We use the CommonCrawl split from Redpajama [71] as $\mathcal{D}$ to exclude the influence of domain weights [80]. During pre-processing, multiple documents may be merged into a single pre-training instance with 1,024 tokens. For the downstream loss $J({\bm{\theta}})$, we compute the LM’s loss on the training split of LIMA [84], a high-quality dataset consisting of 1,030 diverse instruction-response pairs that cover a wide range of downstream scenarios. Our evaluation is conducted on various downstream datasets other than LIMA to avoid over-fitting.
@@
We adopt the same model architecture as Mistral [42] and pre-train LMs with 160M, 470M, 1B, and 1.7B parameters. Model configurations are detailed in Table 6.
@@
To compute the data quality scores from PMP, we adopt a 160M proxy LM. $\mathcal{D}^{\mathrm{prx}}$ consists of 160K instances uniformly sampled from $\mathcal{D}$. We first pre-train the proxy LM on $\mathcal{D}$ for 50K steps and then select checkpoints at $\left[10K,20K,30K,40K,50K\right]$ steps. Initialized from these checkpoints, the proxy LM undergoes inner loops with $\eta=0.008$ over $T^{\mathrm{prx}}=100$ steps with a mini-batch size of 256. ${\bm{\gamma}}^{*}$ is updated for one outer loop epoch with $\alpha=1$. For the data scorer, we fine-tune a 125M Fairseq-Dense model [5] along with the linear head, using the objective in Eq. (8). The training details for the data scorer can be found in Appendix E.2. For Data Selection, we set $\delta=0.1$, $r=0.4$, with further hyper-parameter exploration provided in Appendix G.3.
@@
We pre-train all LMs for 100k steps, using a batch size of 512 and a max input length of 1,024, resulting in roughly 50B tokens. In Section 3.2, we select a 50B-token dataset from a CC corpus containing 125B tokens to assess how different data selection methods improve LM learning given a sufficiently large $\mathcal{D}$. In Section 3.3 (Data-Constrained Setting), we also analyze the effectiveness of PDS when $\mathcal{D}$ is limited in size. See Appendix E.3 for more pre-training details.
@@
We evaluate the LMs’ 0-shot accuracy on the downstream test datasets used in OLMo [31] and their 0/5-shot accuracy on MMLU [39]. We also report the LM’s language modeling loss on a subset of DCLM [48], a high-quality corpus curated with complex pipelines and human heuristics, to verify that models trained on $\mathcal{D}^{\prime}$ retain diversity and long-tail knowledge coverage. Further evaluation details are in Appendix E.4.
@@
  • Conventional: conventionally pre-training LM on 50B tokens uniformly sampled from $\mathcal{D}$, also referring to “Redpajama” in Figure 1.
  • RHO-Loss [58]: selecting data with high reducible losses, as in Eq. (51).
  • DSIR [79]: selecting data with high n-gram overlap with instances in LIMA.
  • IF-Score [43]: selecting data with high influence scores, as in Eq. (52).
@@
Model Size = 160M
Conventional 32.2 34.9 51.4 25.6 40.9 22.5 58.3 65.5 57.6 43.2
RHO-Loss 32.2 35.3 53.2 28.1 40.5 24.1 58.5 63.1 53.0 43.1
DSIR 32.8 35.7 52.5 26.6 41.2 23.8 57.8 68.7 54.7 43.8
IF-Score 32.2 35.7 51.1 27.4 40.8 22.6 57.6 64.1 55.8 43.0
PDS 33.5 38.2 51.4 28.4 42.3 24.1 59.2 68.8 58.7 45.0
@@
Table 1: Results on the downstream evaluation datasets in OLMo [31]. We report the accuracy scores and the average scores across the datasets. The best scores of each model size are boldfaced. PDS achieves the best performance in most cases compared to the baselines.
@@ 3.2 Main Results
We present the evaluation results of the pre-trained LMs on the OLMo evaluation datasets and MMLU in Table 1 and Table 2, respectively. As shown, PDS outperforms the baselines on most datasets, achieving the best overall performance across models with 160M and 470M parameters. Figure 1(b) shows the scaling curves of average accuracy on the OLMo evaluation sets with respect to the model sizes, which range from 160M to 1.7B, demonstrating that the performance improvement remains consistent as the model scales up. These results indicate that the data quality scores obtained in a proxy environment generalize well to various model sizes and downstream tasks.
$N$ Method 0-shot 5-shot
160M Conventional 27.2 26.7
RHO-Loss 27.3 27.1
DSIR 27.4 27.3
IF-Score 27.4 27.1
PDS 27.4 27.7
470M Conventional 27.6 28.5
RHO-Loss 28.4 28.5
DSIR 28.0 28.2
IF-Score 28.4 28.8
PDS 28.9 29.0
Table 2: The 0/5-shot accuracy scores on MMLU. The best scores of each model size are boldfaced.
@@
Besides downstream tasks, we show that PDS also enhances language modeling on a high-quality pre-training corpus. Figure 4 shows the test losses on the DCLM subset for conventionally pre-trained LMs and PDS-trained LMs with 160M, 470M, 1B, and 1.7B parameters. We can see that PDS-trained LMs constantly achieve better performance across various model sizes. In Table 3, we extrapolate the test losses using the Scaling Law [40], showing that the improvements with PDS persist in pre-training recent large LMs, such as GPT-3 [12] and Llama family [72, 73, 21]. Details of the extrapolation are provided in Appendix 7. The DCLM corpus is curated using a complex pipeline based on human heuristics and is verified to be diverse and comprehensive in its coverage of human knowledge. Algorithm 1 offers a principled alternative to the complex pipeline for curating pre-training corpus.
@@
Figure 4: Test losses on the DCLM corpus [48] for 160M, 470M, 1B and 1.7B LMs.
@@
GPT-3 [12] 175B 300B 2.882 2.872
Llama [72] 6.7B 1.0T 2.942 2.896
Llama 2 [73] 70B 2.0T 2.877 2.855
Llama 3.1 [21] 405B 15T 2.851 2.838
Table 3: Test loss extrapolation using the Scaling Law [40]. We predict the test loss when the LM size $N$ and the trained tokens $D$ meet that of GPT-3 175B, Llama 6.7B, Llama 2 70B, and Llama 3.1 405B. The improvements of PDS remain consistent for these LMs.
@@ 3.3 Analysis
Figure 5: Test losses on DCLM corpus [48] in the data-constrained setting. We select data with PDS for different selection ratios $r$ and train the model for multiple epochs to reach the same token number budgets.
@@
In Section 3.2, we assume the original pre-training data is unlimited to ensure that the PDS-selected corpus contains 50B tokens, the same as in conventional pre-training. In practice, when the data before selection is limited, the LM has to be trained on the PDS-selected data for multiple epochs. In Figure 6, we restrict $\mathcal{D}$ to contain 50B tokens and apply PDS with selection ratios $r\in\left[0.125,0.25,0.5\right]$, corresponding to training the LM on the selected data for $\left[8,4,2\right]$ epochs, respectively. We can see that selecting 1/4 data with PDS and training for 4 epochs achieves the lowest test losses, consistent with the findings of Muennighoff et al. [59]. This suggests that PDS improves data utilization as the high-quality web-crawled corpora run out [74]. We extrapolate the loss curve of conventional pre-training with the Scaling Law suggested by [59], and find that it requires another 42B tokens to achieve the performance of PDS ($r=0.25$, 4 Eps), which means PDS reduces the pre-training data requirement by 1.8 times.
@@
In Section 2.3.1, we introduce efficient implementations to solve the data quality scores by running Algorithm 1 for a single outer epoch, leveraging a small proxy LM trained for a limited number of steps. In Figure 6, we use a simulated setting where ${\bm{\gamma}}^{*}$ can be obtained exactly from Algorithm 1 within an affordable computational budget and compare its performance (measured by $J({\bm{\theta}}_{t})$) and overhead (measured in FLOPs) with our efficient implementation. Details of this simulated setting can be found in Appendix E.6. The results show that the efficient implementation significantly reduces computational overhead while preserving most of the performance of the exact solution.
@@
We compare the computational complexity of PDS with pre-training in Table 4. The overhead of running PDS to select data is only about 1/9 of that of pre-training a 1.7B model. More importantly, since PDS is an offline method, the selected corpus can be used to pre-train any number of LMs without additional computational cost. In addition, the offline nature of PDS allows it to seamlessly integrate into recent highly optimized pre-training pipelines [15], requiring only a replacement of the data source without altering the pre-training process.
@@
Table 4: Asymptotic complexity, GPU FLOPs, and actually spent time of different PDS steps and 1.7B model pre-training. $N^{\mathrm{prx}}$ and $D^{\mathrm{prx}}$ are the sizes of the proxy LM and proxy data. $N^{\mathrm{score}}$ is the size of the data scorer. We elaborate on the details of how the complexity and FLOPs are computed in Appendix F. PDS costs little computational overhead compared to pre-training large LMs.
@@ 4 Related Work
Many works have explored data selection approaches to accelerate LM learning or improve downstream performance [4]. Some curate data prior to LM training, which we refer to offline methods, including domain-mixing [80, 25, 26], data pruning [55, 70, 1], sample-wise data selection [78, 20, 79, 35], and data programming [65, 33, 34]. Other works dynamically select data during LM training by adjusting domain-mixing weights [77, 13, 3] or more fine-grained reweighting strategies [24, 30, 36, 68]. This work studies data selection from its general principles, theoretically deriving optimal selection and designing scalable algorithms to implement it.
@@
The principles of optimal control [47] have been shown to be effective in deep learning [7, 53], by treating the forward pass of a multi-layer neural network as a control process where the hidden vectors are state parameters and the model parameters are control variables to optimize. With the help of Pontryagin’s Maximum Principle [63], some works design optimization algorithms with better convergence rates [50, 83] or broader application scenarios [49], and others provide theoretical foundations of neural networks for better interpretation [37, 28]. Unlike these works, we adopt Optimal Control in a novel and “orthogonal” way, by treating the model’s learning pass as the control process.

### v2

@@ 2.3.2 Data Scorer
where $\bm{\phi}$ is the parameters of the data scorer and $\overline{\bm{h}}(\cdot,\cdot)\in\mathbb{R}^{d}$ is the average output hidden states of an LM along the sequence length, with $d$ representing the hidden state size. ${\bm{w}}\in\mathbb{R}^{d},b\in\mathbb{R}$ are the parameters of the linear head. After fine-tuning, we infer the data quality scores of the instances in $\mathcal{D}$ with the data scorer, where the quality score for $x_{n}$ is given by $\gamma(x_{n})={{\bm{w}}^{*}}^{\top}\overline{{\bm{h}}}(x_{n},\bm{\phi^{*}})+b^{*}$.
@@ 2.4 Discussion
Compared to existing offline data selection methods that curate the pre-training corpus before the LM training starts using pattern information (Xie et al., 2023), deduplication (Tirumala et al., 2023; Abbas et al., 2023), or single-step checkpoints (Engstrom et al., 2024), PDS incorporates long-range training dynamics into data selection, as reflected by the “target vector” ${\bm{\lambda}}^{*}_{t}$ in Eq. (5). This can be critical for selecting high-quality instances, given the highly dynamic nature of LM pre-training. Although we run Algorithm 1 in a proxy environment and transfer the quality scores to the large-scale setting via the data scorer, many previous works (Xie et al., 2024; Yu et al., 2024) have shown that data quality information is learnable and transferable across model sizes. Different LMs also share many common training dynamics (Tirumala et al., 2022).
@@
Unlike recent online data selection approaches to incorporate LM training dynamics (Yu et al., 2024; Wang et al., 2024), PDS selects the pre-training corpus offline. This allows PDS to be run only once and used for pre-training multiple LMs of any sizes, without incurring additional computational overhead. The FLOPs needed by PDS in a proxy environment are also negligible compared to the demands of large-scale pre-training, as shown in a complexity analysis in Section 3.3. Besides, the offline nature of PDS makes it flexible to be integrated into optimized pre-training pipelines (Chowdhery et al., 2023) by simply replacing the data sources.
@@ 3.1 Experimental Setup
We use the CommonCrawl split from Redpajama (Together, 2023) as $\mathcal{D}$ to exclude the influence of domain weights (Xie et al., 2024). For the downstream loss $J({\bm{\theta}})$, we compute the LM’s loss on the training split of LIMA (Zhou et al., 2024), a high-quality dataset consisting of 1,030 diverse instruction-response pairs that cover a wide range of downstream scenarios. The results on more choices of $J({\bm{\theta}})$ are shown in Appendix I.5. Our evaluation is conducted on various downstream datasets other than LIMA to avoid over-fitting.
@@
We adopt the same model architecture as Mistral (Jiang et al., 2023) and pre-train LMs with 160M, 470M, 1B, and 1.7B parameters. Model configurations are detailed in Table 6.
@@
To compute the data quality scores from PMP, we adopt a 160M proxy LM. $\mathcal{D}^{\mathrm{prx}}$ consists of 160K instances uniformly sampled from $\mathcal{D}$. We first pre-train the proxy LM on $\mathcal{D}$ for 50K steps and then select checkpoints at $\left[10K,20K,30K,40K,50K\right]$ steps. Initialized from these checkpoints, the proxy LM undergoes inner loops with $\eta=0.008$ over $T^{\mathrm{prx}}=100$ steps with a mini-batch size of 256. ${\bm{\gamma}}^{*}$ is updated for one outer loop epoch with $\alpha=1$. For the data scorer, we fine-tune a 125M Fairseq-Dense model (Artetxe et al., 2022) along with the linear head, using the objective in Eq. (8). The training details for the data scorer can be found in Appendix G.2. For Data Selection, we set $\delta=0.1$, $r=0.4$, with further hyper-parameter exploration provided in Appendix I.5.
@@
We pre-train all LMs for 100k steps, using a batch size of 512 and a max input length of 1,024, resulting in roughly 50B tokens. In Section 3.2, we select a 50B-token dataset from a CC corpus containing 125B tokens to assess how different data selection methods improve LM learning given a sufficiently large $\mathcal{D}$. In Section 3.3 (Data-Constrained Setting), we also analyze the effectiveness of PDS when $\mathcal{D}$ is limited in size. See Appendix G.3 for more pre-training details.
@@
We evaluate the LMs’ 0-shot accuracy on the downstream test datasets used in OLMo (Groeneveld et al., 2024) and their 0-shot performance on MMLU (Hendrycks et al., 2021). We also report the LM’s language modeling loss on a subset of DCLM (Li et al., 2024), a high-quality corpus curated with complex pipelines and human heuristics, to verify that models trained on $\mathcal{D}^{\prime}$ retain diversity and long-tail knowledge coverage. Further evaluation details are in Appendix G.4.
@@
  • Conventional: conventionally pre-training LM on 50B tokens uniformly sampled from $\mathcal{D}$, also refering to “Redpajama” in Figure 1.
  • RHO-Loss (Mindermann et al., 2022): selecting data with high reducible losses, as in Eq. (55).
  • DSIR (Xie et al., 2023): selecting data with high n-gram overlap with instances in LIMA.
  • IF-Score (Koh & Liang, 2017): selecting data with high influence scores, as in Eq. (56).
@@
Model Size = 1B
Conventional 39.9 47.6 52.4 30.6 49.3 26.4 63.1 73.7 60.9 49.3
RHO-Loss 39.8 47.0 53.0 30.8 48.0 26.4 62.9 71.1 61.0 48.9
DSIR 40.8 47.8 53.0 31.2 49.8 26.8 62.7 76.6 58.0 49.6
IF-Score 39.4 47.0 52.6 28.6 49.4 26.4 63.5 74.0 60.5 49.0
PDS 42.1 48.8 54.0 33.4 51.3 28.0 64.1 78.5 58.7 51.0
Table 1: Results on the downstream evaluation datasets in OLMo (Groeneveld et al., 2024). We report the accuracy scores and the average scores across the datasets. The best scores of each model size are boldfaced. PDS achieves the best performance in most cases compared to the baselines.
@@ 3.2 Main Results
$N$ Method 0-shot PPL
470M Conventional 27.6 34.8
RHO-Loss 28.4 33.0
DSIR 28.0 34.0
IF-Score 28.4 31.1
PDS 28.9 27.1
1B Conventional 29.7 26.1
RHO-Loss 30.2 24.9
DSIR 30.0 25.3
IF-Score 30.7 23.6
PDS 31.4 20.5
Table 2: MMLU results. We report the 0-shot accuracy and the perplexity (PPL) on the ground truth answers. The best scores of each model size are boldfaced.
@@
We present the evaluation results of the pre-trained LMs on the OLMo evaluation datasets and MMLU in Table 1 and Table 2, respectively. As shown, PDS outperforms the baselines on most datasets, achieving the best overall performance across models with 470M and 1B parameters. See Appendix I.1 for results on 160M LMs. Figure 1(b) shows the scaling curves of average accuracy on the OLMo evaluation sets with respect to the model sizes, ranging from 160M to 1.7B, demonstrating that the performance improvement remains consistent as the LM scales up. These results indicate that the data quality scores obtained in a proxy environment generalize well to various model sizes and downstream tasks.
@@
Besides downstream tasks, we show that PDS also enhances language modeling on a high-quality pre-training corpus. Figure 4 shows the test losses on the DCLM subset for conventionally pre-trained LMs and PDS-trained LMs with 160M, 470M, 1B, and 1.7B parameters. We can see that PDS-trained LMs constantly achieve better performance across various model sizes. In Table 3, we extrapolate the test losses using the Scaling Law (Hoffmann et al., 2022), showing that the improvements with PDS persist in pre-training recent large LMs, such as GPT-3 (Brown et al., 2020) and Llama family (Touvron et al., 2023a; b; Dubey et al., 2024). Details of the extrapolation are provided in Appendix 9. The DCLM corpus is curated using a complex pipeline based on human heuristics and is verified to be diverse and comprehensive in its coverage of human knowledge. Algorithm 1 offers a principled alternative to the complex pipeline for curating pre-training corpus.
@@
Figure 4: Test losses on the DCLM corpus (Li et al., 2024) for 160M, 470M, 1B and 1.7B LMs.
@@
GPT-3 (Brown et al., 2020) 175B 300B 2.882 2.872
Llama (Touvron et al., 2023a) 6.7B 1.0T 2.942 2.896
Llama 2 (Touvron et al., 2023b) 70B 2.0T 2.877 2.855
Llama 3.1 (Dubey et al., 2024) 405B 15T 2.851 2.838
Table 3: Test loss extrapolation using the Scaling Law (Hoffmann et al., 2022). We predict the test loss when the LM size $N$ and the trained tokens $D$ meet that of GPT-3 175B, Llama 6.7B, Llama 2 70B, and Llama 3.1 405B. The improvements of PDS remain consistent for these LMs.
@@ 3.3 Analysis
Figure 5: Test losses on DCLM corpus (Li et al., 2024) in the data-constrained setting. We select data with PDS for different selection ratios $r$ and train the model for multiple epochs to reach the same token number budgets.
@@
In Section 3.2, we assume the original pre-training data is unlimited to ensure that the PDS-selected corpus contains 50B tokens, the same as in conventional pre-training. In practice, when the data before selection is limited, the LM has to be trained on the PDS-selected data for multiple epochs. In Figure 6, we restrict $\mathcal{D}$ to contain 50B tokens and apply PDS with selection ratios $r\in\left[0.125,0.25,0.5\right]$, corresponding to training the LM on the selected data for $\left[8,4,2\right]$ epochs, respectively. We can see that selecting 1/4 data with PDS and training for 4 epochs achieves the lowest test losses, consistent with the findings of Muennighoff et al. (2023). This suggests that PDS improves data utilization as the high-quality web-crawled corpora run out (Villalobos et al., 2022). We extrapolate the loss curve of conventional pre-training with the Scaling Law suggested by (Muennighoff et al., 2023), and find that it requires another 42B tokens to achieve the performance of PDS ($r=0.25$, 4 Eps), which means PDS reduces the pre-training data requirement by 1.8 times. See Appendix 9 for details of this extrapolation.
@@
In Section 2.3.1, we introduce efficient implementations to solve the data quality scores by running Algorithm 1 for a single outer epoch, leveraging a small proxy LM trained for a limited number of steps. In Figure 6, we use a simulated setting where ${\bm{\gamma}}^{*}$ can be obtained exactly from Algorithm 1 within an affordable computational budget and compare its performance (measured by $J({\bm{\theta}}_{t})$) and overhead (measured in FLOPs) with our efficient implementation. Details of this simulated setting can be found in Appendix G.6. The results show that the efficient implementation significantly reduces computational overhead while preserving most of the performance of the exact solution.
@@
We compare the computational complexity of PDS with pre-training in Table 4. The overhead of running PDS to select data is only about 1/9 of that of pre-training a 1.7B model. More importantly, since PDS is an offline method, the selected corpus can be used to pre-train any number of LMs without additional computational cost. In addition, the offline nature of PDS allows it to seamlessly integrate into recent highly optimized pre-training pipelines (Chowdhery et al., 2023), requiring only a replacement of the data source without altering the pre-training process.
@@
Table 4: Asymptotic complexity, GPU FLOPs, and actually spent time of different PDS steps and 1.7B model pre-training. $N^{\mathrm{prx}}$ and $D^{\mathrm{prx}}$ are the sizes of the proxy LM and proxy data. $N^{\mathrm{score}}$ is the size of the data scorer. We elaborate on the details of how the complexity and FLOPs are computed in Appendix H. PDS costs little computational overhead compared to pre-training large LMs.
@@ 4 Related Work
Many works have explored data selection approaches to accelerate LM learning or improve downstream performance (Albalak et al., 2024). Some curate data prior to LM training, which we refer to offline methods, including domain-mixing (Xie et al., 2024; Fan et al., 2023; Gao et al., 2020), data pruning (Marion et al., 2023; Tirumala et al., 2023; Abbas et al., 2023), sample-wise data selection (Xia et al., 2024b; Du et al., 2023; Xie et al., 2023; Gu et al., 2023), and data programming (Ratner et al., 2016; Gu et al., 2022a; b). Other works dynamically select data during LM training by adjusting domain-mixing weights (Xia et al., 2024a; Chen et al., 2024; Albalak et al., 2023) or more fine-grained reweighting strategies (Fan & Jaggi, 2023; Grangier et al., 2023; Gu et al., 2024; Thakkar et al., 2023). This work studies data selection from its general principles, theoretically deriving optimal selection and designing scalable algorithms to implement it.
@@
The principles of optimal control (Lewis et al., 2012) have been shown to be effective in deep learning (Benning et al., 2019; Liu & Theodorou, 2019), by treating the forward pass of a multi-layer neural network as a control process where the hidden vectors are state parameters and the model parameters are control variables to optimize. With the help of Pontryagin’s Maximum Principle (Pontryagin, 2018), some works design optimization algorithms with better convergence rates (Li et al., 2017; Zhang et al., 2019) or broader application scenarios (Li & Hao, 2018), and others provide theoretical foundations of neural networks for better interpretation (Han et al., 2019; Geshkovski & Zuazua, 2022). Unlike these works, we adopt Optimal Control in a novel and “orthogonal” way, by treating the model’s learning pass as the control process.